Project

General

Profile

NDOS CommissionMtng 20120514

Commissioning Meeting Monday May 14:
Athans, Gavin, Peter ,Peter L., Jon, Jan, Susan, Leon, Martin

Phil Adamson- NUMI Beam Pulses were from tests

NOvA keys for the Neutrino Control room now available
NDSB has been found unlocked several times.
New mailing list: nova_runcoordinator@fnal.govĀ 

Installation:

    New DCMs and TDUs on single timing chain- Every DCM in it's own software timing chain.     Very fast configuration now.

    APD installation- first dozen recovered parylene didn't have have proper shims. 
        5 of the 12 tight APDs working after loosening
        2 more installed properly (1 fail)
        2 new parylene- not working properly
         2 new silicon- working                  

        This week: More recovered parylene (for trial installations), continue installation of         available new coated units.

    TECC and FEBs available. Need a realistic count of what we need to finish NDOS.

DAQ Software: No major changes last week. Automatic error recovery on hold. Prototype data driven trigger on NDOS. BNEV -> SMEM. Implemented simple algorithm. 

    DCMCheckout, Denis requested DCM 4-2   

DATACheck. End time begin time the same at rollover. Email notification will go to data check group and RC. 

DCS GUI cooling send email -> look for help to monitor.